# Current conservation and amplitude regularisation of the Landau problem: Bohm–Madelung description

### 14 Apr 2026 v1

@@ Abstract
The analysis considers two complementary regularisation schemes: a global Fisher–information–based regularisation and a local canonical (shell) Bohm regularisation derived from stationary flux closure. These are applied within distinct classes of stationary flow, characterised by vanishing and nonvanishing component currents.
@@
Regularisation thus acts as a structural reorganisation mechanism in amplitude space, preserving the Landau spectral scale while reorganising the flux-sector structure through branch-wise amplitude–momentum relations, thus establishing branch-wise organisation as a natural framework for describing stationary Bohmian dynamics in the Landau problem.
@@ 1 Introduction
Although the resulting equations are exactly solvable, the quantum potential in the radial sector can exhibit singular behaviour due to its explicit dependence on the amplitude. This motivates the introduction of regularisation schemes to obtain well-behaved solutions.
@@
In recent work [7], regularisation schemes were introduced based on Fisher-information variational principles (global) and canonical closure (local). These render the equations analytically tractable while preserving the spectral structure.
@@
The purpose of the present work is to analyse how this structure manifests in the Landau problem. The analysis is organised around two classifications: (i) stationary current structure and (ii) regularisation scheme. This leads to four regimes, summarised in Fig.  1.
The programme outlined in Fig. 1 is developed in the following sections. Section 2 establishes the Ermakov–Lewis invariant structure. Section 3 analyses nonlinear behaviour in the azimuthal sector. Section 4 develops global and local regularisation and highlights the contrast between sectors. Section 5 constructs admissible branch solutions.
@@ 2.2 Ermakov–Pinney structure and invariants in cylindrical coordinates
The total energy of the system is labeled as $E^{EL}$ and given by
@@ 3 General form of amplitudes for stationary states with nonzero sectorial flux currents
Here, the only condition that imposed on the currents is, $\bf\nabla\cdot J=0$}, with separable current divergence components to have a nonzero contributions. If we use as a separation constant, then component-wise differential equations are written down as
@@
The azimuthal component is thus of primary interest in this work, as it exhibits nonlinearity and coordinate-wise nonlocality induced by the gauge coupling. The stationary current and energy equations for the azimuthal sector are
@@
Hence the angular sector is reduced to the first-order nonlinear system
@@
One can also write Eq. (31) as a $\pi_{\theta}$ differential equation, and the general form is given by
@@
Once $w(\theta)$ is determined, the angular amplitude is recovered as
@@
This representation replaces the second-order ratio $\Theta^{\prime\prime}/\Theta$ with a Riccati-type structure in the logarithmic derivative $w$, which is more suitable for analysing the nonlinear behaviour of the azimuthal sector.
@@ 4 Local regularisation, global extension, and branch structure of the azimuthal sector
In this section, we analyse the stationary system with nonvanishing current components, where the amplitude and momentum equations are nonlinearly coupled at the branch level. This coupling introduces singular-like behaviour that renders the system analytically intractable in its native form. To address this, we first introduce a local regularisation scheme that restores a controlled amplitude–momentum structure and yields analytically tractable solutions in a restricted domain.
A natural question is whether these locally regularised solutions admit a consistent global extension. As shown below, the radial and axial sectors extend globally with modified spectral structure, while the azimuthal sector develops a nonreal-valued amplitude due to gauge-induced coupling. This obstruction necessitates a distinct branch-wise regularisation of the azimuthal amplitude, which is treated separately in Section 5.
@@ 4.1 Global extension of regularised radial and axial sectors
with regular solutions  [13] After a Liouville scaling to reduce $R(r)$ differential equation to a normal form using the transformation $R(r)=\chi_{r}/\sqrt{r}$ and simplification, one obtains a Langer corrected  [12] equation of $\nu$
@@
$\frac{\nu+1}{2}-\frac{\kappa^{2}_{r}}{4\beta}=a_{r}=-n_{r}\quad\text{ $n_{r}$ is nonnegative integer.}$
@@ 4.1.1 Axial sector: global extension
The next section covers the description of azimuthal’s amplitude constrained by zero flux component.
@@ 5 Branch regularisation of the azimuthal amplitude
The solution is restricted to real branches, excluding complex continuations across these singularities, which are not physically admissible under stationary current conditions. The exponential term contains the constant $\kappa$, related to the azimuthal current parameter $C_{\theta}$. Imposing $C_{\theta}\geq 0$ and the stationary constraint $\sum_{i}C_{i}=0$ requires the radial and axial current constants to satisfy $C_{r}+C_{z}\leq 0$. Thus, the admissible angular amplitude is selected by the current branch, and the regularisation in $\theta$ emerges as a consequence of flux-balanced stationary flow rather than an imposed boundary condition.
@@ 6 Physical evolution of branch-regularised amplitudes
Finally, the azimuthal expression at regularisation limit takes the form,
@@
The azimuthal branch develops a singularity at $(2\Phi/\Phi_{0})\theta=1$, which can take continuous values as functions of $B$ and $r$, in this sense, the nontrivial singular structure of the nonzero-current branch is not generated by the canonical regularisation itself, but arises from the nonlocal coupling induced by the mixed gauge term. Finally, $\kappa$ of the exponent terms of $\Theta(\theta)$ carries the same sign of $C_{\theta}$ and for finiteness of the amplitude, it forces $C_{\theta}\geq 0$.
@@ 6.1 Energy spectrum at regularisation limit
As a consequence of branch regularisation, the energy spectrum is determined primarily by the radial sector, which remains analytically well-defined under global extension. The axial contribution retains its free-particle character, while the azimuthal sector, despite its nonlocal and branch-dependent amplitude structure, does not alter the spectral form. The total energy of the system for the shell regularised case Landau problem is identical in its functional form but with a subleading correction through the Langer-type shift $\nu=\sqrt{l^{2}+\tfrac{1}{4}}$, which becomes negligible for large $l$.
$E_{n_{r}\nu}(k_{z})=E_{\perp}(n_{r},\nu)+E_{z}=\frac{\hbar\omega_{c}}{2}\,(2n_{r}+\nu+1)+\frac{\hbar^{2}k_{z}^{2}}{2m},\qquad\omega_{c}=\frac{|e|B}{m},$ (95)
with explicit substitution for $\nu=\sqrt{l^{2}+\tfrac{1}{4}}$ we get
$\boxed{E_{n_{r}l}(k_{z})=E_{\perp}(n_{r},l)+E_{z}=\frac{\hbar\omega_{c}}{2}\,\left(2n_{r}+\sqrt{l^{2}+\tfrac{1}{4}}+1\right)+\frac{\hbar^{2}k_{z}^{2}}{2m}},\qquad\omega_{c}=\frac{|e|B}{m},$ (96)
In particular, the local regularisation or loss of global analyticity in the azimuthal amplitude does not affect the canonical Bohm-regularised spectrum $E^{\mathrm{CBR}}$. The only modification arises from the radial sector through the Langer-type shift $\nu=\sqrt{l^{2}+\tfrac{1}{4}}$.
@@
For identical radial and axial quantum numbers, one finds $E^{{QM}}\leq E^{{EL}}\leq E^{{CBR}}$ for $l\geq 1$. where, $^{QM}, EE$$^{EL}$ and $E^{CBR}$ are energy spectra obtained from standard quantum mechanics, Ermakov-Lewis and canonical Bohm regularisation routes respectively.
@@
The second term lifts the degeneracy with respect to the azimuthal index $l$, which is otherwise present in the standard Landau problem. This contribution arises from the regularised angular structure and may be viewed as a non-cyclotronic correction induced by canonical regularisation.
@@ 7 Conclusion
A consistent canonical Bohm regularisation of a charged particle in a uniform magnetic field leads to a clear structural decomposition of the stationary solution space. The radial and axial sectors admit globally well-defined analytic extensions, while the azimuthal sector, due to gauge-induced coupling, does not support a globally real separable amplitude. This obstruction is resolved through a branch-wise regularisation, which provides a consistent local description of the azimuthal dynamics.
@@
Despite the nonlocal and branch-dependent nature of the azimuthal amplitude, the energy spectrum remains governed by the radial sector and retains the canonical Bohm-regularised form $E^{\mathrm{CBR}}$, with only a Langer-type correction. Thus, the spectral structure remains consistent with standard quantum mechanics in the appropriate limit.
Recent experimental observations of free-electron Landau states have revealed structured current flow patterns that are naturally interpreted within the Bohmian (hydrodynamic) framework [15]. The present formulation provides a complementary theoretical structure in which such flow behaviour emerges from branch-wise organisation under gauge coupling.

### 19 Apr 2026 v2

@@ Abstract
The analysis considers two complementary regularisation schemes: a global Fisher–information–based regularisation and a local canonical (shell) Bohm regularisation derived from stationary flux closure. These are applied within distinct classes of stationary flow, characterised by vanishing and nonvanishing current components.
@@
Regularisation thus acts as a structural reorganisation mechanism in amplitude space, preserving the Landau spectral scale while reorganising the flux-sector structure through branch-wise amplitude–momentum relations, thereby establishing a natural framework for the description of stationary Bohmian dynamics in the Landau problem.
@@ 1 Introduction
Although the resulting equations are exactly solvable, the quantum potential in the radial sector can exhibit singular behaviour due to its explicit dependence on the amplitude. This motivates the introduction of regularisation schemes that yield well-behaved solutions.
@@
In recent work [7], regularisation schemes were introduced based on Fisher-information variational principles (global) and canonical closure (local). These schemes render the equations analytically tractable while preserving the spectral structure.
@@
The purpose of the present work is to analyse how this structure manifests in the Landau problem. The analysis is organised around two classifications: (i) stationary current structure and (ii) regularisation scheme. This leads to a structured classification of solution regimes, summarised in Fig. 1. The primary distinction is set by the stationary current branches ($C_{i}=0$ and $\sum_{i}C_{i}=0,\;C_{i}\neq 0$), which define the underlying solution structure. Within each branch, different regularisation schemes (Ermakov or shell) yield corresponding amplitude constructions.
The programme outlined in Fig. 1 is developed in the following sections. Section 2 establishes the Ermakov–Lewis invariant structure. Section 3 analyses nonlinear behaviour in the azimuthal sector. Section 4 develops global and local regularisation and highlights the contrast between sectors. Section 5 constructs admissible branch solutions. Section 6 discusses representative evolutions of the total amplitudes under different stationary current conditions and their corresponding spectral forms.
@@ 2.2 Ermakov–Pinney structure and invariants in cylindrical coordinates
The total energy of the system is labeled as $E_{\mathrm{EL}}$ and given by
@@ 3 General form of amplitudes for stationary states with nonzero sectorial flux currents
Here, the only condition imposed on the currents is $\bf\nabla\cdot J=0$}, with separable divergence components allowing nonzero sectorial contributions. If we use as a separation constant, then component-wise differential equations are written down as
@@
The azimuthal component is therefore of primary interest in this work, as it exhibits nonlinearity and coordinate-wise nonlocality induced by the gauge coupling. The stationary current and energy equations for the azimuthal sector are
@@
Eqs. (35) and  (36) form the coupled first-order nonlinear system  (37).
@@
Eliminating $w(\theta)$, one obtains a second-order nonlinear differential equation for $\pi_{\theta}(\theta)$, given by
@@
If the first-order system (37) is solved, the logarithmic derivative $w(\theta)$ can be obtained and the angular amplitude reconstructed via
@@
Thus, the amplitude may be recovered either from the $(\pi_{\theta},w)$ system or, equivalently, from the reduced $\pi_{\theta}(\theta)$ equation. In this representation, the second-order ratio $\Theta^{\prime\prime}/\Theta$ is replaced by a Riccati-type structure in $w(\theta)$, which is convenient for analysing the nonlinear behaviour of the azimuthal sector.
@@ 4 Local regularisation, global extension, and branch structure of the azimuthal sector
In this section, we analyse the stationary system with nonvanishing current components, where the amplitude and momentum equations are nonlinearly coupled at the branch level. This coupling introduces near-singular behaviour that renders the system analytically intractable in its native form. To address this, we first introduce a local regularisation scheme that restores a controlled amplitude–momentum structure and yields analytically tractable solutions in a restricted domain.
A natural question is whether these locally regularised solutions admit a consistent global extension. As shown below, the radial and axial sectors extend globally with modified spectral structure, while the azimuthal sector develops a nonreal-valued amplitude due to gauge-induced coupling. This obstruction necessitates a branch-wise regularisation of the azimuthal amplitude, which is treated separately in Section 5.
@@ 4.1 Global extension of regularised radial and axial sectors
with regular solutions  [12]. After a Liouville scaling to reduce the differential for $R(r)$ to a normal form using the transformation $R(r)=\chi_{r}/\sqrt{r}$ and simplification, one obtains a Langer-corrected equation of $\nu$
@@
$\nu+\frac{1}{2}-\frac{\kappa_{r}^{2}}{4\beta}=a_{r}=-n_{r},\qquad n_{r}\text{ is a nonnegative integer.}$
@@ 4.1.1 Axial sector: global extension
The next section examines the azimuthal amplitude under canonical regularisation.
@@ 5 Branch regularisation of the azimuthal amplitude
The solution is restricted to real branches, excluding complex continuations through these singularities, which are not physically admissible under stationary current conditions. The exponential term contains the constant $\kappa$, related to the azimuthal current parameter $C_{\theta}$. Imposing $C_{\theta}\geq 0$ and the stationary constraint $\sum_{i}C_{i}=0$ requires the radial and axial current constants to satisfy $C_{r}+C_{z}\leq 0$. Thus, the admissible angular amplitude is selected by the current branch, and the regularisation in $\theta$ emerges as a consequence of flux-balanced stationary flow rather than an imposed boundary condition.
@@ 6 Physical evolution of branch-regularised amplitudes
Finally, the azimuthal expression at the regularisation limit takes the form,
@@
The azimuthal branch develops a singularity at $(2\Phi/\Phi_{0})\theta=1$, which can take continuous values as functions of $B$ and $r$. In this sense, the nontrivial singular structure of the nonzero-current branch is not generated by the canonical regularisation itself, but arises from the nonlocal coupling induced by the mixed gauge term. Finally, $\kappa$ of the exponent terms of $\Theta(\theta)$ carries the same sign of $C_{\theta}$ and for finiteness of the amplitude, it forces $C_{\theta}\geq 0$.
@@ 6.1 Energy spectrum at regularisation limit
As a consequence of branch regularisation, the energy spectrum is determined primarily by the radial sector, which remains analytically well-defined under global extension. The axial contribution retains its free-particle character, while the azimuthal sector, despite its nonlocal and branch-dependent amplitude structure, does not alter the spectral form. The total energy of the shell regularised Landau problem is identical in its functional form but with a subleading correction through the Langer-type shift $\nu=\sqrt{l^{2}+\tfrac{1}{4}}$, which becomes negligible for large $l$.
$E_{n_{r}\nu}(k_{z})=E_{\perp}(n_{r},\nu)+E_{z}=\frac{\hbar\omega_{c}}{2}\,(2n_{r}+\nu+1)+\frac{\hbar^{2}k_{z}^{2}}{2m},\qquad\omega_{c}=\frac{|e|B}{m}.$ (95)
Upon substituting $\nu=\sqrt{l^{2}+\tfrac{1}{4}}$, one obtains the canonically regularised energy $E_{\mathrm{CBR}}$ as
$\boxed{E_{n_{r}l}(k_{z})=E_{\perp}(n_{r},l)+E_{z}=\frac{\hbar\omega_{c}}{2}\left(2n_{r}+\sqrt{l^{2}+\tfrac{1}{4}}+1\right)+\frac{\hbar^{2}k_{z}^{2}}{2m}}\qquad\omega_{c}=\frac{|e|B}{m}.$ (96)
In particular, the local regularisation or loss of global analyticity in the azimuthal amplitude does not affect the canonical Bohm-regularised spectrum $E_{\mathrm{CBR}}$. The only modification arises from the radial sector through the Langer-type shift $\nu=\sqrt{l^{2}+\tfrac{1}{4}}$.
@@
For identical radial and axial quantum numbers, one finds $E_{\mathrm{QM}}\leq E_{\mathrm{EL}}\leq E_{\mathrm{CBR}}$ for $l\geq 1$. where, $_{\mathrm{QM}}, EE$$_{\mathrm{EL}}$ and $E_{\mathrm{CBR}}$ are energy spectra obtained from standard quantum mechanics, Ermakov-Lewis and canonical Bohm regularisation routes respectively.
@@
The second term lifts the degeneracy with respect to the azimuthal index $l$, which is otherwise present in the standard Landau problem. This contribution arises from the regularised angular structure and may be viewed as a non-cyclotronic contribution induced by canonical regularisation.
@@ 7 Conclusion
A consistent canonical Bohm regularisation of a charged particle in a uniform magnetic field leads to a clear structural decomposition of the stationary solution space. The radial and axial sectors admit globally well-defined analytic extensions, while the azimuthal sector, due to gauge-induced coupling, does not admit a globally real separable amplitude. This obstruction is resolved through a branch-wise regularisation, which provides a consistent local description of the azimuthal dynamics.
@@
Despite the nonlocal and branch-dependent nature of the azimuthal amplitude, the energy spectrum remains governed by the radial sector and retains the canonical Bohm-regularised form $E_{\mathrm{CBR}}$, with only a Langer-type correction. Thus, the spectral structure remains consistent with standard quantum mechanics in the appropriate limit.
Recent experimental observations of free-electron Landau states have revealed structured current flow patterns that are naturally interpreted within the Bohmian (hydrodynamic) framework [14]. The present formulation provides a complementary theoretical structure in which such flow behaviour emerges from branch-wise organisation induced by gauge coupling.